4. Find a basis for the kernel of each of the following linear transformations. 3 -6
Answer #1

Solution:

small 4.left ( a ight )

small varphi _{1}:P_{3} ightarrow mathbb{R}

pi (p) = p (1

small mathrm{Ker}: varphi _{1}=left { pin P_{3}:pleft ( left ( 1 ight )=0 ight ) ight }

small dpi{100} small =left { a+bx+cx^{2}+dx^{3}:a+b+c+d=0 ight }

small =left { a+bx+cx^{2}+dx^{3}:a=-b-c -d ight }

small =left { -b-c-d+bx+cx^{2} +dx^{3} ight }

small =left { bx-b+cx^{2}-c+dx^{3}-d ight }=left { bleft ( x-1 ight )+cleft ( x^{2} -1 ight )+dleft ( x^{3} -1 ight ) ight }

span1( 1), (x2-1), (r -1))

Also, or (r _ 1) + β (12-1)+1(23-1):0

small Rightarrow alpha =eta=gamma =0

small herefore left { left ( x-1 ight ),: : : left ( x^{2} -1 ight ) ,: left ( x^{3} -1 ight ) ight } is linearly independent.

small herefore left { left ( x-1 ight ),: : : left ( x^{2} -1 ight ) ,: left ( x^{3} -1 ight ) ight } is a basis for the kernel.

-------------------------------------------------------------------------------

small left ( b ight )

small varphi _{2}:M_{2 imes 3} ightarrow M_{2 imes 3}

small varphi _{2}left ( B ight )=egin{pmatrix} 3 & -6 -1 &2 end{pmatrix}B

small mathrm{Ker}: varphi _{2}=left { Bin M_{2 imes 3} :varphi _{2}left ( B ight )=egin{pmatrix} 0 & 0 &0 0 & 0 &0 end{pmatrix} ight }

small =left { egin{pmatrix} a & c &e b &d & f end{pmatrix}in M_{2 imes 3} :egin{pmatrix} 3 &-6 -1 & 2 end{pmatrix}egin{pmatrix} a & c &e b& d & f end{pmatrix}=egin{pmatrix} 0 &0 & 0 0 & 0 & 0 end{pmatrix} ight }

small =left { egin{pmatrix} a & c &e b &d & f end{pmatrix}in M_{2 imes 3} :egin{pmatrix} 3a-6b & 3c-6d &3e-6f -a+2b& -c+2d & -e+2f end{pmatrix}=egin{pmatrix} 0 &0 & 0 0 & 0 & 0 end{pmatrix} ight }

small =left { egin{pmatrix} a & c &e b &d & f end{pmatrix}in M_{2 imes 3} :a=2b,: : c=2d,: : e=2f ight }

small =left { egin{pmatrix} 2b& 2d &2f b &d & f end{pmatrix} ight }

0 2 0 00 2 1 0 0

small =mathrm{span}left { egin{pmatrix} 2& 0 &0 1 &0 &0 end{pmatrix} ,egin{pmatrix} 0 & 2 &0 0 & 1 & 0 end{pmatrix},egin{pmatrix} 0 & 0 &2 0&0 & 1 end{pmatrix} ight }

Also, small left { egin{pmatrix} 2& 0 &0 1 &0 &0 end{pmatrix} ,egin{pmatrix} 0 & 2 &0 0 & 1 & 0 end{pmatrix},egin{pmatrix} 0 & 0 &2 0&0 & 1 end{pmatrix} ight }     is linearly independent.

small herefore left { egin{pmatrix} 2& 0 &0 1 &0 &0 end{pmatrix} ,egin{pmatrix} 0 & 2 &0 0 & 1 & 0 end{pmatrix},egin{pmatrix} 0 & 0 &2 0&0 & 1 end{pmatrix} ight } form a basis for the kernel.

We were unable to transcribe this image

We were unable to transcribe this image

pi (p) = p (1

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

span-1), (x2-1)

or (z _ 1) + β (2.2-1)=0

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

0 2 0 00 2 1 0 0

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image